OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-2,-1) and (-5,-3)

OpenStudy (anonymous):

remember the slope formula?

jimthompson5910 (jim_thompson5910):

Answer is \[y=\frac{2}{3}x+\frac{1}{3}\] See attachment for how I got that answer.

OpenStudy (anonymous):

It said incorrect

jimthompson5910 (jim_thompson5910):

what form do they want the equation in?

OpenStudy (anonymous):

y=mx+b using integers or fractions

jimthompson5910 (jim_thompson5910):

how are you entering the answer. Hopefully using an equation editor

OpenStudy (anonymous):

it has one built in

jimthompson5910 (jim_thompson5910):

well what I have is the right answer. I double checked to make sure

jimthompson5910 (jim_thompson5910):

ok and you're using it?

jimthompson5910 (jim_thompson5910):

a lot of the times, the computer is really really picky about how you enter the answer

OpenStudy (anonymous):

no, you are right, I was on the wrong question

jimthompson5910 (jim_thompson5910):

k cool

OpenStudy (anonymous):

{1/6,-1/3} and {5/6,3} thats the one I am on

jimthompson5910 (jim_thompson5910):

Answer is \[y=5x-\frac{7}{6}\] See attached as for how I got that answer.

jimthompson5910 (jim_thompson5910):

Let me know how it works out

jimthompson5910 (jim_thompson5910):

I know this is probably completely overwhelming, but just go over these attachments till you get it. Don't worry, just take it at your pace and you'll do fine.

OpenStudy (anonymous):

How about (-1,-7) and (-6,-4)

jimthompson5910 (jim_thompson5910):

Answer is \[y=-\frac{3}{5}x-\frac{38}{5}\] Again, see image attachment for the work shown.

OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. express your answer in the form y=mx+b (8,9);x+8y=3

jimthompson5910 (jim_thompson5910):

x+8y=C 8+8(9)=C 8+72=C 80=C So the equation is x+8y=80. Now solve for y x+8y=80 8y=-x+80 y=-1/8x+10 Answer: \[y=-\frac{1}{8}x+10\]

OpenStudy (anonymous):

you are the man!

jimthompson5910 (jim_thompson5910):

great, just study over all that we went over and you'll be fine. Glad to have helped out.